1. A 20-cm-diameter disk emits light uniformly from its surface. 40 cm from this disk, along its axis, is a 16.0-cm-diameter opaque black disk; the faces of the two disks are parallel. 40 cm beyond the black disk is a white viewing screen. The lighted disk illuminates the screen, but theres a shadow in the center due to the black disk. What is the diameter of the completely dark part of this shadow?
Answer #1

2 cm 40 cm 10 cm 8 cm 80 cm

In triangle ABC

tanheta = AB/AC

tanheta = 2/40                                                eq-1

In triangle CDE

CD = 8 - r

DE = 40

tanheta = CD/DE

tanheta = (8 - r)/40                                        eq-2

using eq-1 and eq-2

2/40 = (8 - r)/40

r = 6 cm

d = diameter = 2 r = 2 x 6 = 12 cm